# Understanding the Cluster LP for Correlation Clustering

### v2

@@ 6.3 Wrapping up: Deferred proofs of Lemma 37, 38, 39
Another technique that enhances estimation accuracy involves discretizing the $y_{uvw}$ values. We achieve this by subdividing the interval $[max(0,(l(I_{i})+l(I_{j})+l(I_{k})-1)/2),r(I_{k})]$ into numerous sub-intervals. The specific manner in which we partition this interval depends on both the interval’s length and its significance in the overall estimation.
However, it’s crucial to note that discretizing $y_{uvw}$ results in a substantial increase in the number of triangles to be considered. This surge may render the final SDP impractical to solve. The formulation of our ultimate factor-revealing SDP is provided below:
[…]
@@
$\sum_{T\in\mathcal{T_{D}}}\eta_{T}=1,\qquad\qquad\qquad\qquad\qquad\qquad$
@@
For budget functions $b^{+}\equiv b^{+}_{1.437}$ and $b^{-}\equiv b^{-}_{1.437}$ Let ${\mathrm{OPT}_{\mathrm{SDP}}}$ be the optiamal value for factor-revealing SDP, there exist intervals $I_{1},I_{2},...,I_{t}$ and intervals for $y_{uvw}$ such that ${\mathrm{OPT}_{\mathrm{SDP}}}=0$.
@@
For budget functions $b^{+}\equiv b^{+}_{1.437}$ and $b^{-}\equiv b^{-}_{1.437}$, For degenerate and non-degenerate triangles, based on Theorem 59 and Lemma 60, we know that
@@
The graph of the plus edges of our gap instance is based on the line graph of a base graph; given a based graph $H=(V_{H},E_{H})$, our correlation clustering instance is $G=(V_{G},E_{G})$ where $V_{G}=E_{H}$ and $e,f\in V_{G}$ have a plus edge in $G$ if they share a vertex in $V_{H}$.
@@
Let $H=(V_{H},E_{H})$ be a complete graph on $n$ vertices. Let $d=n-1$ be the degree of $H$. Our correlation clustering instance $G=(V_{G},E_{G})$ is the line graph of $H$; $V_{G}=E_{H}$ and $e,f\in E_{H}$ has $+$ edge in $G$ if and only if they share a vertex in $H$. The $+$ degree of each $e\in E_{H}$ in $G$ is $22d-$ 2.
Consider the following solution for the cluster LP: for every $v\in V_{H}$, let $E_{v}\subseteq E_{H}$ be the $d$ edges containing $v$. The cluster LP has $z_{E_{v}}=1/2$ for every $v\in v_{H}$. Each $e\in E_{H}$ belongs to two fractional clusters, each of which has its $d-1$ plus neighbors, so fractionally $d-1$ plus edges incident on it are violated. Since each violated edge is counted twice, the LP value is $\binom{n}{2}(d-1)/2$.
@@
Using this, we can prove a lower bound on the cost of our near-optimal clustering. Note that every cluster is a clique of $+$edges. Thus, the only edges violated are $+$edges. Moreover, there are at most $\sum_{i\in[t]}n_{i}^{2}/2\leq n^{3}/6$ correctly clustered $+$edges. The cost of our near-optimal clustering is the total number of $+$edges of $G$ minus the number of correctly clustered $+$edges, namely at most $\binom{n}{2}(d-1)-n^{3}/6=n^{3}/3-o(n^{3})$. Since the cost of the optimal clustering is at most $35n^{2}$ lower than ours, it is still $n^{3}/3-o(n^{3})$. The fractional solution has the value at most $n^{3}/4$, so the gap is at least $4/3-o(1)$. ∎
@@
In this section, we show that it is NP-hard (under randomized reductions) to obtain an algorithm with an approximation ratio of 24/231.043absent1.043\ $\geq$ 1.043≥ 1.043, proving Theorem 5 restated below.
@@
Still, we use a similar construction for $3$-uniform hypergraphs; let $H=(V_{H},E_{H})$ be the underlying $3$-uniform hypergraph and $G=(V_{G},E_{G})$ be the plus graph of the final Correlation Clustering instance where $V_{G}=E_{H}$ and $e,f\in E_{H}$ has an edge in $G$ if they share a vertex in $H$. We use the following hardness result of Cohen-Addad, Karthik, and Lee [CAL22] that shows that it is hard to distinguish whether $H$ is nearly bipartite or close to a random hypergraph.
@@
  • (YES) If $\phi$ is satisfiable, there exists $U\subseteq V_{H}$ with $|U|=|V_{H}|/2$ that intersects every hyperedge in $E_{H}$. Moreover, for every $u\in U$, |$|\{e\in E_{H}:e\cap U=\{u\}\}|\geq(1/2-\varepsilon)d$.
@@
The same reduction in Theorem 4.1 of (the arXiv version of) [CAL22] yields the desired hardness. In the following, we highlight the difference between the statement of Theorem 4.1 of [CAL22] and our Theorem 63 and briefly explain how our additional properties are satisfied by their reduction.
@@
    Regularity of $H$: Section 4.5 of [CAL22], based on an earlier weighted hard instance, constructs the final hard instance $H=(V_{H},E_{H})$ as a certain random hypergraph where the degree of each vertex $v$ is the sum of independent $\{0,1\}$ variables with the same expected value. This expected value is $\Theta(|V_{H}|^{1.5})$, so the standard Chernoff and union bound argument will show that the degree of each vertex is almost the same with high probability.
@@
    In the (YES) case, for every $u\in U$, |$|\{e\in E_{H}:e\cap U=\{u\}\}|\geq(1/2-\varepsilon)d$: It follows from their construction in Section 4.1. The construction is analogous to Håstad’s celebrated result on Max-3SAT [Hås01] where in the (YES) case, almost three quarters of the clauses have one true literal and almost one quarter have three true literals, so that for each true literal $\ell$, roughly half of the clauses containing $\ell$ has it as the only true literal.
@@
Since two vertices of $H$ have at most $n$ hyperedges containing both of them, let us make at most $n\cdot\binom{10}{2}$ hyperedges in $C$ that contain at least two of $v_{1},\dots,v_{20}$ as singleton clusters; the objective function increases by at most $n\cdot\binom{10}{2}\cdot|C|$. Then partition the remaining $C$ into $E_{1},\dots,E_{20}$ where $E_{i}:=C\cap E_{v_{i}}$. Each $e\in E_{i}$ has at most $2\cdot 20\cdot n$ plus edges in $\cup_{j\neq i}E_{j}$ ($20$ choices for $v_{j}$, $2$ choices for a vertex in $e\ni\{v_{i}\}$, and $n$ choices for hyperedges containing both vertices), so the objective function increases by at most $O(n|C|)$. So, we partitioned $C$ into $C_{1},\dots,C_{r}$ where all the hyperedges in $C_{i}$ share a common endpoint. In total, we increased the objective function by at most $O(n|C|)$. ∎
@@
[…]
@@
[…]
@@
Using this, we can prove a lower bound on the cost of our near-optimal clustering. Note that every cluster is a clique of $+$edges. Thus, the only edges violated are $+$edges. Moreover, there are at most $\sum_{i\in[t]}c_{i}^{2}/2\leq d^{2}n(0.1+O(\sqrt{\varepsilon})$ correctly clustered $+$edges. The cost of our near-optimal clustering is the total number of $+$edges of $G$ minus the number of correctly clustered $+$edges, namely at least $nd^{2}(1/2-0.1-O(\sqrt{\varepsilon}))=nd^{2}(0.4-O(\sqrt{\varepsilon}))$. Since the cost of the optimal clustering is at most $O(n^{2}d)$ lower than ours, it is still $nd^{2}(0.4-O(\sqrt{\varepsilon}))$ using $d=\omega(n)$.
@@
10 Open Problems
We highlight several problems left open by this work.
  • Our algorithm computes a $(1+\epsilon)$-approximate solution for Cluster LP in $n^{\text{poly}(1/\epsilon)}$ time. An intriguing question is whether Cluster LP can be solved in time $f(\epsilon)\cdot\text{poly}(n)$, for some function $f$. (e.g., $O(2^{\text{poly}(1/\epsilon)}n^{100})$).
  • Our analysis of the rounding procedure, which achieves a $1.437$-approximate ratio, currently requires computer assistance. One challenge is to determine whether a similar approximate ratio can be achieved without computer assistance. Additionally, although computer assistance enables us to impose more constraints on $\eta_{T}$—thereby slightly reducing the approximate ratio—it remains substantially above the theoretical lower bound. Exploring the possibility of achieving an approximate ratio of $4/3+\epsilon$, or even surpassing the $4/3$ approximate ratio, represents a critical area for future research.
  • Recent findings by Cohen-Addad, Lolck, Pilipczuk, Thorup, Yan, and Zhang [CALP+24] demonstrate that a 1.847-approximate ratio is achievable in $\tilde{O}(m)$ time, where $m$ is the number of $+$edges. Can we show some kind of tradeoff between the running time and the approximate ratio?
Appendix A Preclustering
The goal of this section is to prove Theorem 14, which is repeated below: See 14
For convenience, we assume every $u$ has a self-loop in $E^{+}$. Let $N^{+}_{u}$ be the set of $+$neighbors of $u$; so we have $u\in N^{+}_{u}$. Let $d^{+}_{u}=|N^{+}_{u}|$ be its $+$degree. Let ${\mathcal{C}}^{*}$ be the optimum clustering, and ${\mathrm{opt}}=\mathrm{obj}({\mathcal{C}}^{*})$ be its cost. We assume $\varepsilon>0$ is at most a small enough constant; we define $\beta=0.1$ to be an absolute constant.
A.1 Constructing Atoms
Our ${\mathcal{K}}$ is obtained from ${\mathcal{C}}$ by marking some vertices and creating singletons for them. We view ${\mathcal{K}}$ both as a clustering and as the set of atoms. The algorithm is described in Algorithm 4.
Algorithm 4 Construction of ${\mathcal{K}}$
1:for every non-singleton $C\in{\mathcal{C}}$ do
5:return ${\mathcal{K}}$
Lemma 66.
Proof.
If a cluster $C$ has less than $\frac{\beta|C|}{3}$ marked vertices before Step 3, then the cost incurred by separating the marked vertices in $C$ is at most $\frac{2}{\beta}=O(1)$ times the cost of edges incident to these vertices in ${\mathcal{C}}$. On the other hand, if $C$ has at least $\frac{\beta|C|}{3}$ marked vertices before Step 3, the cost of edges incident to $C$ in ${\mathcal{C}}$ is at least $\frac{\beta}{2}\cdot|C|\cdot\frac{\beta|C|}{3}\cdot\frac{1}{2}=\Omega(|C|^{2})$. The cost incurred by breaking $C$ into singletons is at most ($|C|\choose 2$. As every edge is charged at most twice, the cost incurred by creating singletons for all marked vertices is at most $O(1)\cdot\mathrm{obj}({\mathcal{C}})$. So $\mathrm{obj}({\mathcal{K}})\leq O(1)\cdot{\mathrm{opt}}$ as $\mathrm{obj}({\mathcal{C}})\leq 3\cdot{\mathrm{opt}}$. ∎
Lemma 67.
For every non-singleton $K\in{\mathcal{K}}$, and every $u\in K$, we have $|N^{+}_{u}\triangle K|<\beta|K|$.
Proof.
Lemma 68.
Proof.
With Lemma 68, we now restrict ourselves to clusterings that do not break atoms. As a result, we can then assume all the edges between two vertices in a same atom $K\in{\mathcal{K}}$ are $+$edges. The assumption only decreases ${\mathrm{opt}}$ and can only make the two properties of Theorem 14 harder to satisfy. We use $K_{u}$ for every $u\in V$ to denote the atom that contains $u$. We let $k_{u}=|K_{u}|$.
It is convenient for us to distribute the $+$edges incident to an atom $K\in{\mathcal{K}}$ equally to the vertices in $K$. For every $u,v\in V$, we define
to be the probability that an edge between a random vertex in $K_{u}$ and a random vertex in $K_{v}$ is a $+$edge. So $w_{uv}\in[0,1]$ for every pair $uv$; in particular, if $v\in K_{u}$, we have $w_{uv}=1$. We shall call $w_{uv}$ the weight of the edge $uv$. For any set $u\in V,V^{\prime}\subseteq V$, we define $w(u,V^{\prime}):=\sum_{v\in V^{\prime}}w_{uv}$ be the total weight of edges between $u$ and $V^{\prime}$. Let $w_{u}:=w(u,V)$ be the total weight of all edges incident to $u$. In this new instance, $w_{uv}$ fraction of edge $uv$ has a $+$ sign, and the remaining $1-w_{uv}$ fraction has a -- sign; $w_{u}$ is the total fractional number of $+$edges incident to $u$, which can be treated as the $+$degree of $u$. This new instance is equivalent to the original one. Till the end of this section we focus on this instance and call it the averaged instance to distinguish it from the original one.
A.2 Defining Admissible Edges
In this section, we give a $(1+O(\varepsilon))$-approximate clustering ${\mathcal{C}}^{*}_{1}$ that does not break atoms. Then we define the set $E_{\mathrm{adm}}$ of admissible edges so that all edges between two different clusters in ${\mathcal{C}}^{*}_{1}$ are admissible. We shall bound $|E_{\mathrm{adm}}|$ in terms of $\mathrm{obj}({\mathcal{K}})$, which is at most $O(1)\cdot{\mathrm{opt}}$. Notice that the construction of ${\mathcal{C}}^{*}_{1}$ is not a part of our algorithm for Theorem 14, but the definition of admissible edges is.
Lemma 69.
Proof.
The inequalilty is by the condition of while loop.
Again, the second inequality is by the condition of the while loop.
We let ${\mathcal{C}}^{*}_{1}$ be the $(1+O(\varepsilon))$-approximate clustering satisfying the properties of Lemma 69. We now proceed to the definition of $E_{\mathrm{adm}}$; this should be independent of ${\mathcal{C}}^{*}_{1}$.
Lemma 70.
Proof.
Then, we define a set $E^{1}$ of edges so that $uv\in E^{1}$ if $\varepsilon w_{v}<w_{u}<\frac{w_{v}}{\varepsilon}$; so $(V,E^{1})$ is undirected and contains a self-loop $(u,u)$ for every $u\in V$. By Lemma 70, we have
Corollary 71.
Also, notice that $E^{1}$ may not be a subset of $E^{+}$. We use $N^{1}_{u}$ to denote the neighbor sets of $u$ in the graph $(V,E^{1})$.
Lemma 72.
Proof.
The second inequality is by that $w_{vp},w_{up}\in[0,1]$. ∎
Define a set $E^{2}$ of edges so that $uv\in E^{2}$ if and only if $v\in K_{u}$, or the following holds: $uv\in E^{1}$ and $\sum_{p\in N^{1}_{u}\cap N^{1}_{v}}w_{up}w_{vp}>\varepsilon(w_{u}+w_{v})$. Notice that $E^{2}\subseteq E^{1}$. Corollary 71 and Lemma 72 implies
Corollary 73.
Let $d^{2}_{u}$ and $N^{2}_{u}$ denote the degree and neighbor set of $u$ in the graph $(V,E^{2})$ respectively.
Lemma 74.
$d^{2}_{u}-k_{u}\leq O(\frac{1}{\varepsilon^{2}})(w_{u}-k_{u})$.
Proof.
Therefore, we can define $uv$ to be admissible, i.e., in $E_{\mathrm{adm}}$, if $uv\in E^{2}$ and $v\notin K_{u}$. This finishes the description of the preclustered instance $({\mathcal{K}},E_{\mathrm{adm}})$. The total number of admissible edges is
Appendix B Detail of Lemma 60
We repeat the factor-revealing SDP for convenience.
[…]
$\sum_{T\in\mathcal{T_{D}}}\eta_{T}=1,\qquad\qquad\qquad\qquad\qquad\qquad$
We aim to demonstrate that ${\mathrm{OPT}_{\mathrm{SDP}}}=0$ for the budget functions $b^{+}\equiv b^{+}_{1.437}$ and $b^{-}\equiv b^{-}_{1.437}$, where ${\mathrm{OPT}_{\mathrm{SDP}}}$ is the optimal value for factor-revealing SDP. We describe the specific interval choices and small tricks employed to expedite the computation of factor-revealing SDP.
We opt to divide the interval $[0,1]$ into 37 sub-intervals with the following splitting points:
[0.0,0.1[0.0,0.1,[ 0.0 , 0.1 , $0.15,0.2,0.3,0.32,1/3,0.34,0.36,0.38,0.4,0.42,0.43,0.44,0.45,0.46,0.47,0.48,0.49,0.5,0.51,$
0.52,0.53,0.54,0.56,0.58,0.6,0.62,0.65,0.7,0.8,0.9,0.95,0.96,0.97,0.98,0.99,1 0.52,0.53,0.54,0.56,0.58,0.6,0.62,0.65,0.7,0.8,0.9,0.95,0.96,0.9% 7,0.98,0.99,1]0.52 , 0.53 , 0.54 , 0.56 , 0.58 , 0.6 , 0.62 , 0.65 , 0.7 , 0.8 , 0.9 , 0.95 , 0.96 , 0.97 , 0.98 , 0.99 , 1 ]
For intervals $I_{i},I_{j},I_{k}$ where $I_{i}\in[0.42,0.60]$ and $I_{j}\in[0.42,0.60]$, we discretize the $y_{uvw}$ value. For each $I_{i},I_{j},I_{k}$, we split the interval
$I_{l}=[max(0,(l(I_{i})+l(I_{j})+l(I_{k})-1)/2),r(I_{k})]$
into 10 sub-intervals, each with the same length. For triangles in different ranges but with the same $y_{uv},y_{uw},y_{vw}$ value, they are treated as different triangles since they affect different locations in the $Q$ matrix. However, if two triangles have the same $y_{uv},y_{uw},y_{vw},y_{uvw}$ and are located in the same $I_{i},I_{j},I_{k}$, but are located in different $I_{l}$ intervals, they can be treated as the same triangle since they affect $Q$ in the same way. We may need to use different $\tilde{d}$ values to lower bound $T$ when $T$ is in a different range. Assume $I^{(\tau)}_{l}$ and $I^{(\tau+1)}_{l}$ share $y_{uvw}=r(I^{(\tau)}_{l})=l(I^{(\tau+1)}_{l})$ value. To lower bound $\Delta(T)-\text{cost}(T)$ in different ranges, we set $\tilde{d}(T)=\min_{T_{1}\in\mathcal{T}(I_{i},I_{j},I_{k},I^{(\tau)}_{l})\cup\mathcal{T}(I_{i},I_{j},I_{k},I^{(\tau+1)}_{l})}\Delta(T_{1})-\text{cost}(T_{1})$. This trick reduces the number of triangles to consider when discretizing $y_{uvw}$, considering only 12 triangles instead of 20.
We choose to create numerous subintervals, especially around 0.5, to obtain precise estimates for $+$\!+\!-+ + - triangles with $y$ values of $(y_{uv}=0.5,y_{uw}=0.5,y_{vw}=0,y_{uvw}=0)$ and ++++\!+\!++ + + triangles with $y$ values of $(y_{uv}=0.5,y_{uw}=0.5,y_{vw}=0.5,y_{uvw}=0.5)$. When determining the intervals, our strategy involves setting up a budget function and identifying instances where ${\mathrm{OPT}_{\mathrm{SDP}}}$ becomes negative, allowing us to refine the interval splitting for these problematic cases.
Approximately 274,000 triangles are considered for the current interval configuration. The SDP solver is executed on Google Cloud with N2 CPUs, taking around 44,000 seconds to produce the ${\mathrm{OPT}_{\mathrm{SDP}}}$. It’s important to note that loading all triangles into memory poses a bottleneck, requiring approximately 200 GB of memory.
Lower bound of the factor-revealing SDP
One might wonder whether using more intervals and denser splitting of $y_{uvw}$’s interval could demonstrate ${\mathrm{OPT}_{\mathrm{SDP}}}\geq 0$ for a smaller budget function. Unfortunately, even with more subintervals, a significant improvement is not achievable if we only use $Q$ and $F$ as constraints.
Lemma 75.
For the budget functions $b^{+}\equiv b^{+}_{1.421}$ and $b^{-}\equiv b^{-}_{1.421}$, there does not exist intervals $I_{1},I_{2},...,I_{t}$ and intervals for $y_{uvw}$ such that ${\mathrm{OPT}_{\mathrm{SDP}}}\geq 0$.
Proof.
Consider the $+$\!+\!-+ + - triangle $T_{1}$ with $y$ value $(y_{uv}=0.5,y_{uw}=0.5,y_{vw}=0,y_{uvw}=0)$ and ++++\!+\!++ + + triangles $T_{2}$ with $y$ values of $(y_{uv}=0.5,y_{uw}=0.5,y_{vw}=0.5,y_{uvw}=0.5)$. If $\eta_{T_{1}}=0.75$ and $\eta_{T_{2}}=0.25$, both $Q$ and $F$ are PSD. Additionally, $\Delta(T_{1})-\text{cost}(T_{1})\approx-0.1819$ and $\Delta(T_{2})-\text{cost}(T_{2})\approx 0.54519$, leading to ${\mathrm{OPT}_{\mathrm{SDP}}}\leq\eta_{T_{1}}(\Delta(T_{1})-\text{cost}(T_{1}))+\eta_{T_{2}}(\Delta(T_{2})-\text{cost}(T_{2}))<0$. ∎
Appendix C Standard SDP Relaxation

### v3

@@ 4.3 Using Independently Sampled Clusters to Construct Cluster LP Solution
Another technique that enhances estimation accuracy involves discretizing the $y_{uvw}$ values. We achieve this by subdividing the interval $max(0,(l(I_{i})+l(I_{j})+l(I_{k})-1)/2),r(I_{k})]$ into numerous subintervals. The specific manner in which we partition this interval depends on both the interval’s length and its significance in the overall estimation.
However, it is crucial to note that discretizing $y_{uvw}$ results in a substantial increase in the number of triangles to be considered. This surge may render the final SDP impractical to solve. The formulation of our ultimate factor-revealing SDP is provided below.
$\min\qquad\sum_{I_{i},I_{j},I_{k},I_{l}}\sum_{T\in\mathcal{T_{D}}(I_{i},I_{j},I_{k},I_{l})}\eta_{T}\cdot\tilde{d}(I_{i},I_{j},I_{k},I_{l})\qquad\text{s.t.}$ (factor-revealing SDP)
@@
$\sum_{T\in\mathcal{T_{D}}}\eta_{T}=1.\qquad\qquad\qquad\qquad\qquad\qquad$
@@
For Algorithm 4, let $\mathrm{budget}^{+}\equiv\mathrm{budget}^{+}_{1.485}$ and $\mathrm{budget}^{-}\equiv\mathrm{budget}^{-}_{1.485}$. Let $\mathrm{OPT}_{\mathrm{SDP}}$} be the optimum of the factor-revealing SDP factor-revealing SDP below. There exists a discretization of the edge and triple variables into intervals $I_{1},\ldots,I_{t}$ (together with an interval grid for $y_{uvw}$) such that ${\mathrm{OPT}_{\mathrm{SDP}}}\geq 0$.
@@
For budget functions $\mathrm{budget}^{+}\equiv\mathrm{budget}^{+}_{1.485}$ and $\mathrm{budget}^{-}\equiv\mathrm{budget}^{-}_{1.485}$, and for degenerate and non-degenerate triangles, based on Theorem 51 and Lemma 52, we know that
@@
The graph of the ++edges of our gap instance is based on the line graph of a base graph; given a base graph $H=(V_{H},E_{H})$, our Correlation Clustering instance is $G=(V_{G},E_{G})$ where $V_{G}=E_{H}$ and $e,f\in V_{G}$ are connected by a ++edge in $G$ if they share a vertex in $V_{H}$.
@@
Let $H=(V_{H},E_{H})$ be a complete graph on $n$ vertices. Let d=n1d=n-1 be the degree of $H$. Our Correlation Clustering instance $G=(V_{G},E_{G})$ is the line graph of $H$; $V_{G}=E_{H}$ and $e,f\in E_{H}$ has ++ edge in $G$ if and only if they share a vertex in $H$. The ++ degree of each $e\in E_{H}$ in $G$ is 2d22d-2.
Consider the following solution for the cluster LP: for every $v\in V_{H}$, let $E_{v}\subseteq E_{H}$ be the $d$ edges containing $v$. The cluster LP has $z_{E_{v}}=1/2$ for every $v\in v_{H}$. Each $e\in E_{H}$ belongs to two fractional clusters, each of which has its d1d-1 plus neighbors, so fractionally d1d-1 plus edges incident on it are violated. Since each violated edge is counted twice, the LP value is $\binom{n}{2}(d-1)/2$.
@@
Using this, we can prove a lower bound on the cost of our near-optimal clustering. Note that every cluster is a clique of ++edges. Thus, the only edges violated are ++edges. Moreover, there are at most $\sum_{i\in[t]}n_{i}^{2}/2\leq n^{3}/6$ correctly clustered ++edges. The cost of our near-optimal clustering is the total number of ++edges of $G$ minus the number of correctly clustered ++edges, namely at most $\binom{n}{2}(d-1)-n^{3}/6=n^{3}/3-o(n^{3})$. Since the cost of the optimal clustering is at most $35n^{2}$ lower than ours, it is still $n^{3}/3-o(n^{3})$. The fractional solution has the value at most $n^{3}/4$, so the gap is at least 4/3o(1)4/3-o(1). ∎
@@
In this section, we show that it is NP-hard (under randomized reductions) to obtain an algorithm with an approximation ratio of 24/231.043\ $\geq$ 1.043, proving Theorem 5 restated below.
@@
We still use a similar construction for $3$-uniform hypergraphs; let $H=(V_{H},E_{H})$ be the underlying $3$-uniform hypergraph and $G=(V_{G},E_{G})$ be the plus graph of the final Correlation Clustering instance where $V_{G}=E_{H}$ and $e,f\in E_{H}$ has an edge in $G$ if they share a vertex in $H$. We use the following hardness result of Cohen-Addad, Karthik, and Lee [CKL22] that shows that it is hard to distinguish whether $H$ is nearly bipartite or close to a random hypergraph.
@@
  • (YES) If $\phi$ is satisfiable, there exists $U\subseteq V_{H}$ with $|U|=|V_{H}|/2$ that intersects every hyperedge in $E_{H}$. Moreover, for every $u\in U$, $|\{e\in E_{H}:e\cap U=\{u\}\}|\geq(1/2-\varepsilon)d$.
@@
The same reduction in Theorem 4.1 of (the arXiv version of) [CKL22] yields the desired hardness. In the following, we highlight the difference between the statement of Theorem 4.1 of [CKL22] and our Theorem 55 and briefly explain how our additional properties are satisfied by their reduction.
@@
    Regularity of $H$: Section 4.5 of [CKL22], based on an earlier weighted hard instance, constructs the final hard instance $H=(V_{H},E_{H})$ as a certain random hypergraph where the degree of each vertex $v$ is the sum of independent $\{0,1\}$ variables with the same expected value. This expected value is $\Theta(|V_{H}|^{1.5})$, so the standard Chernoff and union bound argument will show that the degree of each vertex is almost the same with high probability.
@@
    In the (YES) case, for every $u\in U$, $|\{e\in E_{H}:e\cap U=\{u\}\}|\geq(1/2-\varepsilon)d$: It follows from their construction in Section 4.1. The construction is analogous to Håstad’s celebrated result on Max-3SAT [Hås01] where in the (YES) case, almost three quarters of the clauses have one true literal and almost one quarter have three true literals, so that for each true literal $\ell$, roughly half of the clauses containing $\ell$ has it as the only true literal.
@@
Since two vertices of $H$ have at most $n$ hyperedges containing both of them, let us make at most $n\cdot\binom{10}{2}$ hyperedges in $C$ that contain at least two of $v_{1},\dots,v_{20}$ as singleton clusters; the objective function increases by at most $n\cdot\binom{10}{2}\cdot|C|$. Then partition the remaining $C$ into $E_{1},\dots,E_{20}$ where $E_{i}:=C\cap E_{v_{i}}$. Each $e\in E_{i}$ has at most $2\cdot 20\cdot n$ plus edges in $\cup_{j\neq i}E_{j}$ (2020 choices for $v_{j}$, $2$ choices for a vertex in $e\ni\{v_{i}\}$, and $n$ choices for hyperedges containing both vertices), so the objective function increases by at most $O(n|C|)$. So, we partitioned $C$ into $C_{1},\dots,C_{r}$ where all the hyperedges in $C_{i}$ share a common endpoint. In total, we increased the objective function by at most $O(n|C|)$. ∎
@@
(Note that $|E_{H}|\leq(1+\delta)dn/3$.) Interpreting $\frac{1}{n}\sum_{j=1}^{i}f_{j/n}$ as $\int_{0}^{1}f(x)dx$ where $f(x)=c_{\lceil xn\rceil}$, we have that $\sum_{i=1}^{t}|c_{i}|^{2}\leq(1+\delta)^{2}d^{2}n\max_{f}\int_{0}^{1}f(x)^{2}dx$, where the maximum is taken over functions $f:[0,1]\to[0,1]$ with the constraints that
@@
Then one see that the optimal $f$ satisfies either $f(y)=1$ or $\int_{x=0}^{y}f(x)=y-y^{2}+y^{3}+\varepsilon/3$ for every $y\in[0,1)$. If it is not satisfied at some $y$, we can increase $f(y)$ while decreasing $f(z)$ for some $z>y$, which will still satisfy the constraints and increase $\int_{0}^{1}f(x)^{2}dx$. Therefore, we can conclude that $f(y)=1$ for $y\leq\tau$ and
$\int_{x=0}^{y}f(x)dx=y-y^{2}+y^{3}/3+\varepsilon/3\Rightarrow f(y)=(y-y^{2}+y^{3}/3+\varepsilon/3)^{\prime}=1-2y+y^{2}$
@@
Using this, we can prove a lower bound on the cost of our near-optimal clustering. Note that every cluster is a clique of ++edges. Thus, the only edges violated are ++edges. Moreover, there are at most $\sum_{i\in[t]}c_{i}^{2}/2\leq d^{2}n(0.1+O(\sqrt{\varepsilon})$ correctly clustered ++edges. The cost of our near-optimal clustering is the total number of ++edges of $G$ minus the number of correctly clustered ++edges, namely at least $nd^{2}(1/2-0.1-O(\sqrt{\varepsilon}))=nd^{2}(0.4-O(\sqrt{\varepsilon}))$. Since the cost of the optimal clustering is at most $O(n^{2}d)$ lower than ours, it is still $nd^{2}(0.4-O(\sqrt{\varepsilon}))$ using $d=\omega(n)$.
@@
Appendix A Standard SDP Relaxation